what would you add to 4x^2-6x to make it a complete squre? a.4/9 b.9/16 c.16/25 d.4/25 e.9/25
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

what would you add to 4x^2-6x to make it a complete square?
---------
If you square ax-b you get:
a^2x^2 - 2abx + b^2
-----
Your problem:
a^2 = 4, so a = 2.
-2ab = -6, so -2*2*b = -6 ; So b = -6/-4 = 3/2
---------------------------
You need to add b^2 = 9/4
---
None of the answers you listed is correct.
Cheers,
Stan H.



a.4/9 b.9/16 c.16/25 d.4/25 e.9/25

Answer by checkley77(12844)

4x^2-6x+2.25
(2x-1.5)^2 ans.
Proof:
2x-1.5
2x-1.5 multiply.
------------
4x^2-3x-3x+2.25
4x^2-6x+2.25
None of these answers=2.25.
9/4=2.25 ??????
----------------------------
a.4/9
b.9/16
c.16/25
d.4/25
e.9/25